Express the confidence interval 157.2 < μ < 260.6 in the form of ¯ x ± M E .
Answer #1

ANSWER:-157.22M< < 26016 X-M. E <McK+ME 2 Compare both the equations. R-ME = 167.2 a TIME Ž 260.6 Add eq @ and 2 = 260.6+15712 T = 4